OpenStudy (dylanfwlr):

Find the distance between points R(-4,1) and S(2,7)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (dylanfwlr):

very much

hartnn (hartnn):

so what distance did u get ?

OpenStudy (dylanfwlr):

working it out

OpenStudy (dylanfwlr):

i got square root of -72.?

OpenStudy (solomonzelman):

You can't get a negative distance, redo it.

hartnn (hartnn):

-72 ? or +72 ?

OpenStudy (dylanfwlr):

ill say +72 ;)

hartnn (hartnn):

yeah, sqrt 72 is correct :)

hartnn (hartnn):

you can simplify that further if you want

OpenStudy (dylanfwlr):

that's not an answer choice..

hartnn (hartnn):

because that can be simplified

OpenStudy (dylanfwlr):

how?

hartnn (hartnn):

can u factor 72 ?

OpenStudy (dylanfwlr):

1,2,3,4,6,8,9,12,18,24,36,72?

hartnn (hartnn):

nice! so, 72 = 2*2*2*3*3 agreed ?

OpenStudy (dylanfwlr):

yeah

hartnn (hartnn):

\(\sqrt {72} = \sqrt {2^2*3^2 *2}\)

hartnn (hartnn):

\(\sqrt {72 }= 2*3\sqrt 2\)

OpenStudy (dylanfwlr):

thanks lots

hartnn (hartnn):

welcome ^_^